OpenStudy (anonymous):

part a-graph the quadratic linear system) y=-x squared-4x+12 and y=-2x+4 -i understand how to do that. part b) find all the solutions to the system. <---- i dont understand this, please help!

OpenStudy (anonymous):

solution to the system geometrically represent the point of intersection you can get the point of intersectn. by simultaneously solving the equation of curves here, y= -x^2-4x+12 and, y=-2x+4 -x^2-4x+12=-2x+4